OpenStudy (anonymous):

Long division polynomial Integral not sure what long division is going on here .. probably just forgetting something

OpenStudy (anonymous):

\[\int \frac{dx}{x^2-x^3+\frac{5}{12}x^4-\frac{7}{72}x^5+....}\]

OpenStudy (anonymous):

it goes to then next step saying long after long division

OpenStudy (anonymous):

the way my teacher did it was he had the outside the bottom and the inside 1... but that just seems weird to me

OpenStudy (unklerhaukus):

i can't even see the pattern in the denominator yet, by i guess you can factor out \(x^2\)

OpenStudy (anonymous):

still not seeing what long division is going lol sorry i know it's going to be